OpenStudy (anonymous):

Please help Meh Which of the following options is an equivalent function to f(x) = 3(2)^3^x? (both 3 and x are listed as exponents :)

OpenStudy (anonymous):

please help @phi

OpenStudy (phi):

is it \[ f(x) = 3(2)^{3^x} \]?

OpenStudy (phi):

or \[ f(x)= 3(2)^{3x} \] ?

OpenStudy (phi):

we can't do much with the first. the second can be written as \[ f(x)= 3(2^3)^x = 3\cdot 8^x\]

OpenStudy (anonymous):

whats the difference between the two equations? @phi sorry my mum called me

OpenStudy (phi):

the first one has \(3^x\) as the exponent. the second way has 3x as the exponent. the first way can't be simplified. the second way can be, using the rule \[ a^{bc}= (a^b)^c \]

OpenStudy (anonymous):

(3x) is an exponent, they are both together and the same location/size... @phi

OpenStudy (phi):

that means you can write the 2^(3x) as either \[ (2^x)^3 \] or \[ (2^3)^x\] the second way means \[ (2\cdot 2\cdot 2)^x \] or \[8^x\]

OpenStudy (phi):

and we still have a 3 out front so you could write the expression 3 * 2^(3x) as 3*8^x

OpenStudy (phi):

Does that match any of your options?